Given the following information for octane g/mol), the total amount of absorbed 540 octane is heated from -20°C to 1450C. Boiling Point 125.7 °C 56.8 °C Melting Point kJ Vap 41.5 mol kJ AHFusion 20.7 mol cp (l) 254.6 moi °C cp (g) 316.9 mol °C
Answer #1

Since melting point - 56.8 0C and we are asked to calculate from 20 degree C, it indicates that the compound is in liquid state.

q1 = m * s * (t2-t1)lquid

q1 = Heat change occured from - 20 to 125.7 degree C=(540/114.6)*254.6*(125.7+20) = 174794. J

q2 = heat of vaporisation = (540/114.6) * 41500 = 195550. J

q3 = Heat change occured from 125.7 to 145 degreeC = (540/114.6) * 316.9 * (145 - 125.7) = 28820.J

Total heat change = q1 + q2 + q3 = 399164 J = 399.kJ